# ReconX: Reconstruct Any Scene from Sparse Views with Video Diffusion Model

### v2

@@ 4.1 Overview of ReconX
Given $K$ sparse-view (i.e., as few as two) images =$\mathcal{I}=\left\{\bm{I}^{i}\right\}_{i=1}^{K},\left(\bm{I}^{i}\in\mathbb{R}^{H\times W\times 3}\right)$, our goal is to reconstruct the underlying 3D scene, where we can synthesize novel views of unseen viewpoints. In our framework ReconX, we first build a global point cloud $\mathcal{P}=\left\{\bm{p}_{i},1\leq i\leq N\right\}\in\mathbb{R}^{N\times 3}$ from $\mathcal{I}$ and project $\mathcal{P}$ into the 3D context representation space $\mathcal{F}$ as the structure condition $\mathcal{F}(\mathcal{P})$ (Sec. 4.2). Then we inject $\mathcal{F}(\mathcal{P})$ into the video diffusion process to generate 3D consistent video frames =$\mathcal{I}^{\prime}=\left\{\bm{I}^{i}\right\}_{i=1}^{K^{\prime}},(K^{\prime}>K)$, thus creating more observations (Sec. 4.3). To alleviate the negative artifacts caused by the inconsistency among generated videos, we utilize the confidence maps $\mathcal{C}=\left\{\mathcal{C}_{i}\right\}_{i=1}^{K^{\prime}}$ from the DUSt3R model and LPIPS loss (Zhang et al., 2018a) to achieve a robust 3D reconstruction (Sec. 4.4). In this way, we can unleash the full power of the video diffusion model to reconstruct intricate 3D scenes from very sparse views. Our pipeline is depicted in Figure 2.
4.2 Building the 3D Structure Condition
@@
Grounded by the theoretical analysis in Sec. 3, we leverage an unconstrained stereo 3D reconstruction method DUSt3R (Wang et al., 2024a) with point-based representations to build the 3D structure condition $\mathcal{F}$. Given a set of sparse images =$\mathcal{I}=\left\{\bm{I}^{i}\right\}_{i=1}^{K}$, we first construct a connectivity graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ of $K$ input views similar to DUSt3R, where vertices $\mathcal{V}$ and each edge $e=(n,m)\in\mathcal{E}$ indicates that the images $\bm{I}^{n}$ and $\bm{I}^{m}$ shares visual contents. Then we use $\mathcal{G}$ to recover a globally aligned point cloud $\mathcal{P}$. For each image pair $e=(n,m)$, we predict pairwise pointmaps ${P}^{n,n},{P}^{m,n}$ and their corresponding confidence maps $\mathcal{C}^{n,n},\mathcal{C}^{m,n}\in\mathbb{R}^{H\times W\times 3}$. For clarity, we denote ${P}^{n,e}:={P}^{n,n}$ and ${P}^{m,e}:=\mathcal{P}^{m,n}$. Since we aim to rotate all pairwise predictions into a shared coordinate frame, we introduce transformation matrix $T_{e}$ and scaling factor $\sigma_{e}$ associated with each pair $e\in\mathcal{E}$ to optimize global point cloud $\mathcal{P}$ as:
@@
More details of the point cloud extraction can be found in Wang et al. (2024a). Having aligned the point clouds $\mathcal{P}$, we now project it into a 3D context representation space $\mathcal{F}$ through a transformer-based encoder for better interaction with latent features of the video diffusion model. Specifically, we embed the input point cloud $\mathcal{P}$ into a latent code using a learnable embedding function and a cross-attention encoding module:
@@
where $\tilde{\mathcal{P}}$ is a down-sampled version of $\mathcal{P}$ at $1/8$ scale to efficiently distill input points to a compact 3D context space. Finally, we get the 3D structure guidance $\mathcal{F}(\mathcal{P})$ which contains sparse structural information of the 3D scene that can be interpreted by the denoising U-Net.
4.3 3D Consistent Video Frames Generation
In this subsection, we incorporate the 3D structure condition $\mathcal{F}(\mathcal{P})$ into the video diffusion process to obtain 3D consistent frames. To achieve consistency between generated frames and high-fidelity rendering views of the scene, we utilize the video interpolation capability to recover more unseen observations, where the first frame and the last frame of input to the video diffusion model are two reference views. Specifically, given sparse-view images =$\mathcal{I}=\left\{\bm{I}^{i}_{\text{ref}}\right\}_{i=1}^{K}$ as input, we aim to render consistent frames $f(\bm{I}^{i-1}_{\text{ref}},\bm{I}^{i}_{\text{ref}})=\{\bm{I}^{i-1}_{\text{ref}},\bm{I}_{2},...,\bm{I}_{T},\bm{I}^{i}_{\text{ref}}\}\in\mathbb{R}^{(T+2)\times 3\times H\times W}$ where $T$ is the number of generated novel frames. To unify the notation, we denote the embedding of image condition in the pretrained video diffusion model as $F_{g}=g(\bm{I_{\text{ref}}})$ and the embedding of 3D structure condition as $F_{\mathcal{F}}=\mathcal{F}(\mathcal{P})$. Subsequently, we inject the 3D condition into the video diffusion process by interacting with the U-Net intermediate feature $F_{\text{in}}$ through the cross-attention of spatial layers:
@@
where $Q=F_{\text{in}}W_{Q},K_{g}=F_{g}W_{K},V_{g}=F_{g}W_{V}$, $K_{\mathcal{F}}=F_{\mathcal{F}}W_{K}^{\prime},V_{\mathcal{F}}=F_{\mathcal{F}}W_{V}^{\prime}$ are the query, key, and value of 2D and 3D embeddings respectively. $W_{Q},W_{K},W_{K}^{\prime},W_{V},W_{V}^{\prime}$ are the projection matrices and $\lambda_{\mathcal{F}}$ denotes the coefficient that balances image-conditioned and 3D structure-conditioned features. Given the first and last two views condition $c_{\text{view}}$ from $F_{g}$ and 3D structure condition $c_{\text{struc}}$ from $F_{\mathcal{F}}$, we apply the classifier-free guidance (Ho & Salimans, 2022) strategy to incorporate the condition and our training objective is:
@@
$\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{\bm{x}\sim p,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\bm{x}_{t},t,c_{\text{view}},c_{\text{struc}}\right)\right\|_{2}^{2}\right],$ (5)
@@
where $\bm{x}_{t}$ is the noise latent from the ground-truth views of the training data.
4.4 Confidence-Aware 3DGS Optimization.
Built upon the well-designed 3D structure condition, our video diffusion model generates highly consistent video frames, which can be used to reconstruct the 3D scene. As conventional 3D reconstruction methods are originally designed to handle real-captured photographs with calibrated camera metrics, directly applying these approaches to the generated videos is not effective to recover the coherent scene due to the uncertainty of unconstrained images (Wang et al., 2024a; Fan et al., 2024). To alleviate the uncertainty issue, we adopt a confidence-aware 3DGS mechanism to reconstruct the intricate scene. Different from recent approaches (Martin-Brualla et al., 2021a; Ren et al., 2024) which model the uncertainty in per-image, we instead focus on a global alignment among a series of frames. For the generated frames $\left\{\bm{I}^{i}\right\}_{i=1}^{K^{\prime}}$, we denote $\hat{C}_{i}$ and $C_{i}$ as the per-pixel color value for predicted and generated view $i$. Then, we model the pixel values as a Gaussian distribution in our 3DGS, where the mean and variance of $\bm{I}^{i}$ are $C_{i}$ and $\sigma_{i}$. The variance $\sigma_{i}$ measures the discrepancy between the predicted and generated images. The uncertainty metric $\sigma_{i}$ for each image is estimated by minimizing the following negative log-likelihood among all frames:
$\mathcal{L}_{I_{i}}=-\log\left(\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{\|\hat{C}_{i}-C_{i}^{{}^{\prime}}\|_{2}^{2}}{2\sigma^{2}}\right)\right).$ (6)
where ${C}_{i}^{{}^{\prime}}=\mathcal{A}({C}_{i},\{{C}_{i}\}_{i=1}^{K^{\prime}}\setminus{C}_{i})$ and $\mathcal{A}$ is a tailored global align function to establish connections between each frame and the other frames, enabling a more robust global uncertainty estimation. Specifically, the training objective of DUSt3R is to map image pairs to 3D space, while the confidence map $\mathcal{C}$ represents the model’s confidence in the pixel matches of image pairs within the 3D scene. Through its training process, DUSt3R inherently assigns low confidence to mismatched regions in image pairs, achieving the goal of Eq. 6. The confidence maps $\left\{\mathcal{C}_{i}\right\}_{i=1}^{K^{\prime}}$ for each generated frames $\left\{\bm{I}^{i}\right\}_{i=1}^{K^{\prime}}$ are equivalent to the uncertainty $\sigma_{i}$. Meanwhile, the pairwise matching between all frames accomplishes the global alignment operation $\mathcal{A}$. Moreover, we introduce the LPIPS (Zhang et al., 2018b) loss to remove the artifacts and further enhance the visual quality. Towards this end, we formulate the confidence-aware 3DGS loss between the Gaussian rendered image $\hat{\bm{I}}^{i}$ and generated frame ${\bm{I}}^{i}$ as:
$\mathcal{L}_{\text{conf}}=\sum_{i=1}^{K^{\prime}}\mathcal{C}_{i}\left(\lambda_{\text{rgb}}\mathcal{L}_{1}(\hat{\bm{I}}^{i},\bm{I}^{i})+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}}(\hat{\bm{I}}^{i},\bm{I}^{i})+\lambda_{\text{lpips}}\mathcal{L}_{\text{lpips}}(\hat{\bm{I}}^{i},\bm{I}^{i})\right).$ (7)
5 Experiments
In this section, we conduct extensive experiments to evaluate our sparse-view reconstruction framework ReconX. We first present the setup of the experiment (Sec 5.1). Then we report our qualitative and quantitative results compared to feed-forward based methods (Sec 5.2) and per-scene optimization-based methods (Sec 5.3) in various settings. Finally, we conduct ablation studies to further verify the efficacy of our framework design (Sec 5.4). Please refer to our supplementary materials for more comparisons and visualizations.
5.1 Experiment Setup
Implementation Details. In our framework, we choose DUSt3R (Wang et al., 2024a) as our unconstrained stereo 3D reconstruction backbone and the I2V model DynamiCrafter (Xing et al., 2023) (@ $512\times 512$ resolution) as the video diffusion backbone. We first finetune the image cross-attention layers with 2000 steps on the learning rate $1\times 10^{-4}$ for warm-up. Then we incorporate the 3D structure condition $c_{\text{struc}}$ into the video diffusion model and further finetune the spatial layers with 30K steps on the learning rate of $1\times 10^{-5}$. Our video diffusion was trained on 3D scene datasets by sampling 32 frames with dynamic FPS at the resolution of $512\times 512$ in a batch. The AdamW (Loshchilov & Hutter, 2017) optimizer is employed for optimization. At the inference of our video diffusion, we adopt the DDIM sampler (Song et al., 2022) using multi-condition classifier free guidance (Ho & Salimans, 2022). Similar to Xing et al. (2023), we adopt tanh gating to learn $\lambda_{\mathcal{F}}$ adaptively. The training is conducted on 8 NVIDIA A800 (80G) GPUs in two days. In the 3DGS optimization stage, we choose the point maps of the first and end frames as the initial global point cloud and all 32 generated frames are used to reconstruct the scene. Our implementation follows the pipeline of the original 3DGS (Kerbl et al., 2023), but unlike this method, we omit the adaptive control process and attain high-quality renderings in just 1000 steps. The coefficients $\lambda_{\text{rgb}}$, $\lambda_{\text{ssim}}$, and $\lambda_{\text{lpips}}$ are set to 0.8, 0.2, and 0.5, respectively.
Datasets. The video diffusion model of ReconX is trained on three datasets: RealEstate-10K (Zhou et al., 2018), ACID (Liu et al., 2021), and DL3DV-10K (Ling et al., 2024) based on the pretrained model. RealEstate-10K is a dataset downloaded from YouTube, which is split into 67,477 training scenes and 7,289 test scenes. The ACID dataset consists of natural landscape scenes, with 11,075 training scenes and 1,972 testing scenes. DL3DV-10K is a large-scale outdoor dataset containing 10,510 videos with consistent capture standards. For each scene video, we randomly sample 32 contiguous frames with random skips and serve the first and last frames as the input for our video diffusion model. To further validate our strong generalizability, we also directly evaluate our method on the DTU (Jensen et al., 2014), NeRF-LLFF (Mildenhall et al., 2019), and more challenging outdoor datasets Mip-NeRF 360 (Barron et al., 2022) and Tank-and-Temples dataset (Knapitsch et al., 2017).
Baselines and Metrics. To comprehensively demonstrate our strong capability in sparse-view reconstruction, we compare our ReconX with (a) feed-forward based methods trained from 3D scenes to learn 3D prior and (b) per-scene optimization based methods with specific priors (e.g., , depth) for sparse-view reconstruction. Specifically, we compare with NeRF-based pixelNeRF (Yu et al., 2021) and MuRF (Xu et al., 2024); Light Field based GPNR (Suhail et al., 2022) and AttnRend (Du et al., 2023); and the recent state-of-the-art 3DGS-based pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024a) in feed-forward based comparisons. On the other hand, we compare with SparseNeRF (Wang et al., 2023), original 3DGS (Kerbl et al., 2023), and DNGaussian (Li et al., 2024) for per-scene optimization comparisons. Furthermore, we qualitatively compare our method with more recent works CAT3D (Gao et al., 2024) and ReconFusion (Wu et al., 2024b) that incorporate generative power. For quantitative results, we report the standard metrics in NVS, including PSNR, SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018b).
@@ 5.1 Experiment Setup
Figure 3: Qualitative comparison with feed-forward based methods. We provide the comparison of our ReconX with other baselines in Easy Set, Hard Set, and Cross Set. In comparison to these feed-forward based methods, ReconX achieves better visual quality and generalization.
5.2 Comparison with Feed-Forward based Baselines
Comparison for small angle variance in input views. For fair comparison with baseline methods like MuNeRF (Xu et al., 2024), pixelSplat (Charatan et al., 2024), and MVSplat (Chen et al., 2024a), we first compare our reconX with baseline method from sparse views with small angle variance (see Easy Set from Table 2 and Figure 3). We observe that our ReconX surpasses all previous state-of-the-art models in terms of all metrics on visual quality and qualitative perception.
Comparison for large angle variance in input views. As MVSplat and pixelSplat are much better than previous baselines, we conduct thorough comparisons with them in more difficult settings. In more challenging settings (i.e., given sparse views with large angle variance), our proposed ReconX demonstrate more significant improvement than baselines, especially in unseen and generalized viewpoints (see Hard Set from Table 2 and Figure 3). This clearly shows the effectiveness of ReconX in creating more consistent observations from video diffusion to mitigate the inherent ill-posed sparse-view reconstruction problem.
Cross-dataset generalization. Unleashing the strong generative power of the video diffusion model through 3D structure condition, our ReconX is inherently superior in generalizing to out-of-distribution novel scenes. To demonstrate the strong generalizability of ReconX, we conduct two cross-dataset evaluations. For a fair comparison, we train the models solely on the RealEstate10K and directly test them on two popular NVS datasets (i.e., NeRF-LLFF (Mildenhall et al., 2019) and DTU (Jensen et al., 2014)). As shown in Cross Set from Table 2 and Figure 3, the competitive baseline methods MVSplat (Chen et al., 2024a) and pixelSplat (Charatan et al., 2024) fail to render such OOD datasets which contain different camera distributions and image appearance, leading to dramatic performance degradation. In contrast, our ReconX shows impressive generalizability and the gain is larger when the domain gap from training and test data becomes larger.
Assessing more-view quality. ReconX is agnostic to the number of input views. Specifically, given $N$ views as input, we sample a plausible camera trajectory to render image pairs using our video diffusion models and finally optimize the 3D scene from all generated frames. For a fair comparison with Chen et al. (2024a), we verify this by testing on DTU with three context views. Our results are PSNR: 22.83, SSIM: 0.512, LPIPS: 0.317, MVSplat’s are PSNR: 14.30, SSIM: 0.508, LPIPS: 0.371, and pixelSplat’s are PSNR: 12.52, SSIM: 0.367, LPIPS: 0.585. Compared to the two-view results (Table 2), our ReconX and MVSplat both achieve better performance given more input views while we are much better than MVSplat. However, pixelSplat performs worse when using more views also shown in Chen et al. (2024a).
@@ 5.2 Comparison with Feed-Forward based Baselines
Table 1: Quantitative comparisons with feed-forward based methods for small angle variance (Easy Set) in input views. For each scene, the model takes two views as input and renders three novel views for evaluation.
@@
pixelSplat 16.83 0.476 0.494 19.62 0.730 0.270
MVSplat 16.49 0.466 0.486 19.97 0.732 0.245
ReconX 24.53 0.847 0.083 23.70 0.867 0.143
Table 2: Quantitative comparison with feed-forward based methods for large angle variance (Hard Set) in input views and cross-dataset (Cross Set) comparisons to evaluate generalization ability.

### v3

@@ IV-B Overview of ReconX
Given $K$ sparse-view (i.e., as few as two) images =$\mathcal{I}=\left\{\boldsymbol{I}^{i}\right\}_{i=1}^{K},\left(\boldsymbol{I}^{i}\in\mathbb{R}^{H\times W\times 3}\right)$, our goal is to reconstruct the underlying 3D scene, where we can synthesize novel views of unseen viewpoints. In our framework ReconX, we first build a global point cloud $\mathcal{P}=\left\{\boldsymbol{p}_{i},1\leq i\leq N\right\}\in\mathbb{R}^{N\times 3}$ from $\mathcal{I}$ and project $\mathcal{P}$ into the 3D context representation space $\mathcal{F}$ as the structure condition $\mathcal{F}(\mathcal{P})$ (Sec. IV-C). Then we inject $\mathcal{F}(\mathcal{P})$ into the video diffusion process to generate 3D consistent video frames =$\mathcal{I}^{\prime}=\left\{\boldsymbol{I}^{i}\right\}_{i=1}^{K^{\prime}},(K^{\prime}>K)$, thus creating more observations (Sec. IV-D). To alleviate the negative artifacts caused by the inconsistency among generated videos, we utilize the confidence maps $\mathcal{C}=\left\{\mathcal{C}_{i}\right\}_{i=1}^{K^{\prime}}$ from the DUSt3R model and LPIPS loss [54] to achieve a robust 3D reconstruction (Sec. IV-E). In this way, we can unleash the full power of the video diffusion model to reconstruct intricate 3D scenes from very sparse views. Our pipeline is depicted in Fig. 2.
IV-C Building the 3D Structure Condition
@@
Grounded by the theoretical analysis in Sec. IV-A, we leverage an unconstrained stereo 3D reconstruction method DUSt3R [23] with point-based representations to build the 3D structure condition $\mathcal{F}$. Given a set of sparse images =$\mathcal{I}=\left\{\boldsymbol{I}^{i}\right\}_{i=1}^{K}$, we first construct a connectivity graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ of $K$ input views similar to DUSt3R, where vertices $\mathcal{V}$ and each edge $e=(n,m)\in\mathcal{E}$ indicates that the images $\boldsymbol{I}^{n}$ and $\boldsymbol{I}^{m}$ shares visual contents. Then we use $\mathcal{G}$ to recover a globally aligned point cloud $\mathcal{P}$. For each image pair $e=(n,m)$, we predict pairwise pointmaps ${P}^{n,n},{P}^{m,n}$ and their corresponding confidence maps $\mathcal{C}^{n,n},\mathcal{C}^{m,n}\in\mathbb{R}^{H\times W\times 3}$. For clarity, we denote ${P}^{n,e}:={P}^{n,n}$ and ${P}^{m,e}:=\mathcal{P}^{m,n}$. Since we aim to rotate all pairwise predictions into a shared coordinate frame, we introduce transformation matrix $T_{e}$ and scaling factor $\sigma_{e}$ associated with each pair $e\in\mathcal{E}$ to optimize global point cloud $\mathcal{P}$ as:
@@
More details of the point cloud extraction can be found in [23]. Having aligned the point clouds $\mathcal{P}$, we now project it into a 3D context representation space $\mathcal{F}$ through a transformer-based encoder for better interaction with latent features of the video diffusion model. Specifically, we embed the input point cloud $\mathcal{P}$ into a latent code using a learnable embedding function and a cross-attention encoding module:
@@
where $\tilde{\mathcal{P}}$ is a down-sampled version of $\mathcal{P}$ at $1/8$ scale to efficiently distill input points to a compact 3D context space. Finally, we get the 3D structure guidance $\mathcal{F}(\mathcal{P})$ which contains sparse structural information of the 3D scene that can be interpreted by the denoising U-Net. The PosEmb is a column-wise positional embedding function: $\mathbb{R}^{3}\rightarrow\mathbb{R}^{C}$, where $C$ is the dimension of embedding. More specifically, the PosEmb function is implemented as follows: (1) Fixed Sinusoidal Basis: The basis $\mathbf{e}$ is a 3D sinusoidal encoding: $\mathbf{e}=[\sin(2^{0}\pi p),\sin(2^{1}\pi p),\dots]$, where $p\in\mathbb{R}^{3}$ is the position. (2) Embedding Calculation: The input $\mathbf{x}$ is projected onto $\mathbf{e}$ and its sine and cosine are concatenated: $\mathbf{embeddings}=\text{concat}(\sin(\mathbf{proj}),\cos(\mathbf{proj}))$. (3) Learnable Transformation: The positional encoding is passed through an MLP along with the input $\mathbf{x}$: $\mathbf{y}=\text{MLP}(\text{concat}(\mathbf{embeddings},\mathbf{x}))$. In short, PosEmb combines a fixed sinusoidal encoding with a learnable MLP transformation.
For the transformer-based encoder, we encode the DUSt3R point cloud data to a fixed-length sparse representation of the point cloud. Specifically, we first employ a subsampling based on farthest point sampling (FPS) to reduce the point cloud to a smaller set of key points while retaining its overall structural characteristics. Then, we apply cross-attention between the embeddings of the original point cloud and downsampled point cloud. This mechanism can be interpreted as a form of partial self attention, where the downsampled points act as query anchors that aggregate information from the original point cloud. The encoder is not initialized from any pretrained models. Instead, it is trained jointly with the video diffusion model in an end-to-end manner. This design choice ensures that the encoder is specifically adapted to the characteristics of DUSt3R point clouds in our experiment datasets.
IV-D 3D Consistent Video Frames Generation
In this subsection, we incorporate the 3D structure condition $\mathcal{F}(\mathcal{P})$ into the video diffusion process to obtain 3D consistent frames. To achieve consistency between generated frames and high-fidelity rendering views of the scene, we utilize the video interpolation capability to recover more unseen observations, where the first frame and the last frame of input to the video diffusion model are two reference views. Specifically, given sparse-view images =$\mathcal{I}=\left\{\boldsymbol{I}^{i}_{\text{ref}}\right\}_{i=1}^{K}$ as input, we aim to render consistent frames $f(\boldsymbol{I}^{i-1}_{\text{ref}},\boldsymbol{I}^{i}_{\text{ref}})=\{\boldsymbol{I}^{i-1}_{\text{ref}},\boldsymbol{I}_{2},...,\boldsymbol{I}_{T},\boldsymbol{I}^{i}_{\text{ref}}\}\in\mathbb{R}^{(T+2)\times 3\times H\times W}$ where $T$ is the number of generated novel frames. To unify the notation, we denote the embedding of image condition in the pretrained video diffusion model as $F_{g}=g(\boldsymbol{I_{\text{ref}}})$ and the embedding of 3D structure condition as $F_{\mathcal{F}}=\mathcal{F}(\mathcal{P})$. Subsequently, we inject the 3D condition into the video diffusion process by interacting with the U-Net intermediate feature $F_{\text{in}}$ through the cross-attention of spatial layers:
@@
where $Q=F_{\text{in}}W_{Q},K_{g}=F_{g}W_{K},V_{g}=F_{g}W_{V}$, $K_{\mathcal{F}}=F_{\mathcal{F}}W_{K}^{\prime},V_{\mathcal{F}}=F_{\mathcal{F}}W_{V}^{\prime}$ are the query, key, and value of 2D and 3D embeddings respectively. $W_{Q},W_{K},W_{K}^{\prime},W_{V},W_{V}^{\prime}$ are the projection matrices and $\lambda_{\mathcal{F}}$ denotes the coefficient that balances image-conditioned and 3D structure-conditioned features. Given the first and last two views condition $c_{\text{view}}$ from $F_{g}$ and 3D structure condition $c_{\text{struc}}$ from $F_{\mathcal{F}}$, we apply the classifier-free guidance [55] strategy to incorporate the condition and our training objective is:
@@
$\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{\boldsymbol{x}\sim p,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\boldsymbol{x}_{t},t,c_{\text{view}},c_{\text{struc}}\right)\right\|_{2}^{2}\right],$ (12)
@@
where $\boldsymbol{x}_{t}$ is the noise latent from the ground-truth views of the training data.
@@
Figure 3: Qualitative comparison with two-view 3D scene reconstruction. We provide the comparison with other baselines in Easy Set, Hard Set, and Cross Set. In comparison to these two-views novel view synthesis methods, ReconX achieves better visual quality and generalization.
TABLE I: Quantitative comparisons with feed-forward based methods for small angle variance (Easy Set) in input views. For each scene, the model takes two views as input and renders three novel views for evaluation.
@@
TABLE II: Quantitative comparison with feed-forward based methods for large angle variance (Hard Set) in input views and cross-dataset (Cross Set) comparisons to evaluate generalization ability.
@@
pixelSplat 11.42 0.312 0.611 12.89 0.382 0.560
MVSplat 11.60 0.353 0.425 13.94 0.473 0.385
ReconX 21.05 0.768 0.178 19.78 0.476 0.378
IV-E Confidence-Aware 3DGS Optimization
Built upon the well-designed 3D structure condition, our video diffusion model generates highly consistent video frames, which can be used to reconstruct the 3D scene. As conventional 3D reconstruction methods are originally designed to handle real-captured photographs with calibrated camera metrics, directly applying these approaches to the generated videos is not effective to recover the coherent scene due to the uncertainty of unconstrained images [23, 24]. To alleviate the uncertainty issue, we adopt a confidence-aware 3DGS mechanism to reconstruct the intricate scene. Different from recent approaches [56, 57] which model the uncertainty in per-image, we instead focus on a global alignment among a series of frames. For the generated frames $\left\{\boldsymbol{I}^{i}\right\}_{i=1}^{K^{\prime}}$, we denote $\hat{C}_{i}$ and $C_{i}$ as the per-pixel color value for predicted and generated view $i$. Then, we model the pixel values as a Gaussian distribution in our 3DGS, where the mean and variance of $\boldsymbol{I}^{i}$ are $C_{i}$ and $\sigma_{i}$. The variance $\sigma_{i}$ measures the discrepancy between the predicted and generated images. The uncertainty metric $\sigma_{i}$ for each image is estimated by minimizing the following negative log-likelihood among all frames:
$\mathcal{L}_{I_{i}}=-\log\left(\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{\|\hat{C}_{i}-C_{i}^{{}^{\prime}}\|_{2}^{2}}{2\sigma^{2}}\right)\right),$ (13)
where ${C}_{i}^{{}^{\prime}}=\mathcal{A}({C}_{i},\{{C}_{i}\}_{i=1}^{K^{\prime}}\setminus{C}_{i})$ and $\mathcal{A}$ is a tailored global align function to establish connections between each frame and the other frames, enabling a more robust global uncertainty estimation. Specifically, the training objective of DUSt3R is to map image pairs to 3D space, while the confidence map $\mathcal{C}$ represents the model’s confidence in the pixel matches of image pairs within the 3D scene. Through its training process, DUSt3R inherently assigns low confidence to mismatched regions in image pairs, achieving the goal of Eq. 13. The confidence maps $\left\{\mathcal{C}_{i}\right\}_{i=1}^{K^{\prime}}$ for each generated frames $\left\{\boldsymbol{I}^{i}\right\}_{i=1}^{K^{\prime}}$ are equivalent to the uncertainty $\sigma_{i}$. Meanwhile, the pairwise matching between all frames accomplishes the global alignment operation $\mathcal{A}$. Moreover, we introduce the LPIPS [58] loss to remove the artifacts and further enhance the visual quality. Towards this end, we formulate the confidence-aware 3DGS loss between the Gaussian rendered image $\hat{\boldsymbol{I}}^{i}$ and generated frame ${\boldsymbol{I}}^{i}$ as:
$\mathcal{L}_{\text{conf}}$ $=\sum_{i=1}^{K^{\prime}}\mathcal{C}_{i}\left(\lambda_{\text{rgb}}\mathcal{L}_{1}(\hat{\boldsymbol{I}}^{i},\boldsymbol{I}^{i})+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}}(\hat{\boldsymbol{I}}^{i},\boldsymbol{I}^{i})\right.$
$\quad\left.+\lambda_{\text{lpips}}\mathcal{L}_{\text{lpips}}(\hat{\boldsymbol{I}}^{i},\boldsymbol{I}^{i})\right),$ (14)
V Experiments
In this section, we conduct extensive experiments to evaluate our ReconX. We first present the setup of the experiment (Sec V-A). Then we report our qualitative and quantitative results compared to 3D scene reconstruction method from two-views (Sec V-B) and multi-views (Sec. V-C) in various settings. We also show more generalizable experiment results to evaluate of extrapolation ability (Sec. V-D). Finally, we conduct ablation studies to further verify the efficacy of our framework design (Sec. V-E).
V-A Experiment Setup
Implementation Details. In our framework, we choose DUSt3R [23] as our unconstrained stereo 3D reconstruction backbone and the I2V model DynamiCrafter [18] (@ $512\times 512$ resolution) as the video diffusion backbone. We first finetune the image cross-attention layers with 2000 steps on the learning rate $1\times 10^{-4}$ for warm-up. Then we incorporate the 3D structure condition $c_{\text{struc}}$ into the video diffusion model and further finetune the spatial layers with 30K steps on the learning rate of $1\times 10^{-5}$. Our video diffusion was trained on 3D scene datasets by sampling 32 frames with dynamic FPS at the resolution of $512\times 512$ in a batch. The AdamW [59] optimizer is employed for optimization. At the inference of our video diffusion, we adopt the DDIM sampler [60] using multi-condition classifier free guidance [55]. Similar to [18], we adopt tanh gating to learn $\lambda_{\mathcal{F}}$ adaptively. The training is conducted on 8 NVIDIA A800 (80G) GPUs in two days. In the 3DGS optimization stage, we choose the point maps of the first and end frames as the initial global point cloud and all 32 generated frames are used to reconstruct the scene. Our implementation follows the pipeline of the original 3DGS [2], but unlike this method, we omit the adaptive control process and attain high-quality renderings in just 1000 steps. The coefficients $\lambda_{\text{rgb}}$, $\lambda_{\text{ssim}}$, and $\lambda_{\text{lpips}}$ are set to 0.8, 0.2, and 0.5, respectively.
Figure 4: Qualitative comparison with sparse-view reconstruction methods on Mip-Nerf 360 and Tank and Temples. With sparse views as input, our ReconX achieves much better reconstruction quality compared with baselines.
TABLE III: Quantitative comparisons with multi-views reconstruction methods on MipNeRF 360  and Tank and Temples, and DL3DV. We evaluate the reconstruction performance with different input views for each scene.